# Superluminal Wave Activation at Relativistic Magnetized Shocks

### v1

@@ Abstract
Fast radio bursts (FRBs) are extremely energetic radio transients, some are generated in magnetar magnetospheres and winds. Despite a growing number of observations, their emission mechanisms remain elusive. It has recently been proposed that Alfvénic perturbations can convert into superluminal O-modes at magnetized shocks and propagate in the downstream as a radio signal. We validate this superluminal wave activation mechanism using pair-plasma theory and particle-in-cell simulations. Theory predicts two different downstream modes: non-propagating Alfvénic perturbations and propagating superluminal O-modes. Superluminal wave activation occurs if the frequency of upstream perturbations in the shock frame exceeds the downstream plasma frequency. 1D particle-in-cell simulations confirm wavenumber and frequency jumps across the shock for upstream perturbations with frequencies well above the plasma frequency. Our simulations model both monochromatic upstream waves and broadband spectra with the downstream plasma frequency acting like a high-pass filter for superluminal O-modes. We discuss implications for FRB generation in relativistic magnetized winds.
@@ I Introduction
Relativistic magnetized shocks generate many observable phenomena in extreme astrophysical environments, like particle acceleration to ultra-high energies (e.g., Sironi2009; Sironi2015) and coherent wave generation via synchrotron maser emission (Iwamoto2017; Plotnikov2019; Babul2020; Sironi2021; Vanthieghem2025). Magnetars are neutron stars with strong magnetic fields $B\gtrsim 10^{14}\,{\rm G}$, slow rotation periods $P\lesssim 1\,{\rm s}$, and coronae of magnetically active plasmas (Rea2010; Kaspi2017). Their strongly magnetized and dynamic environments are a natural place for the formation of relativistic shocks at various scales (e.g., Margalit2018; Metzger2019; Beloborodov2020; Thompson2022). Bursting, likely caused by crustal activity, can drive magnetospheric instabilities and seed magnetic fluctuations in expanding relativistic outflows.
@@
Thompson2022 suggested that non-propagating Alfvénic perturbations advected towards a magnetized shock convert to propagating superluminal kinetic plasma waves downstream. Such superluminal O-modes have frequencies $\omega/\bar{\omega}_{\rm p}\gtrsim 1$, where $\bar{\omega}_{\rm p}$ is the rest frame plasma frequency (Arons1986; Bransgrove2023). Superluminal O-modes likely escape compact object environments as radio emission.
@@
Transient radio activity, especially fast radio bursts (FRBs), are messengers of magnetically active environments both within (SGR 1935+2154; Andersen2020) and outside of our Galaxy (frbcollaboration2025). These powerful millisecond-duration flares are typically narrow band (a few hundred MHz) and their peak frequency ranges between $120\,{\rm MHz}$ and several GHz (PastorMarazuela2021). They can carry information about their host environment, for example in polarization (e.g., niu2024), and scintillation (nimmo2024). FRBs are emerging as powerful probes of turbulent plasmas throughout the universe (connor2025; Ocker2025). However, their generation mechanisms remain elusive. This letter provides testable predictions of the Thompson2022 model for radio wave generation at relativistic shocks in compact object magnetospheres. It is organized as follows. Section II outlines a simplified theoretical framework with relevant scales of the wave activation process. Section III shows 1D PIC simulations (setup in Section III.1) of mode activation for monochromatic upstream fluctuations (III.2) and broadband seed wave spectra (III.3). We discuss results and their astrophysical implications in Section IV.
II 1D Model of Alfvén Wave Activation
@@
Figure 1: Schematic visualization of mode conversion at relativistic magnetized shocks. A shock front propagates through a magnetized plasma. Non-propagating Alfvénic perturbations upstream of the shock ($\omega_{A}=0$, waves in white) convert into a superposition of non-propagating Alfvénic perturbations (white) and propagating superluminal O-modes (blue) in the downstream. Quantities measured in different frames are denoted as follows: downstream (d), upstream (u), shock (s).
@@
We consider a 1D magnetized planar shock that propagates along $\hat{\mathbf{x}}$ in a pair plasma with background magnetic field $\bm{B}_{0\rm u}]^{\rm d}=\left(B_{0\rm u}/B_{0}\right)\hat{\mathbf{y}}$ (see Figure 1). Here, $eB_{0}=m_{e}\omega_{Le0}c$ normalizes the magnetic field to the electron Larmor frequency $\omega_{Le0}$ in the downstream frame. Subscripts denote to which region the quantity belongs: the downstream (d), shock-front (s), or upstream (u). Superscripts indicate in which frame a quantity is measured. This section reviews ordinary modes in cold 1D plasmas and describes a mode-conversion mechanism between subluminal (Alfvén) and superluminal modes at magnetized shocks (Figure 1). Non-propagating Alfvénic perturbations frozen into the upstream flow are activated as superluminal modes propagating into the downstream.
II.1 1D Ordinary Modes in Cold Plasmas
@@
In 1D and with $\bm{B}_{0\rm u}$ perpendicular to $\hat{\mathbf{x}}$, only modes with wavenumber $k_{\parallel}=0$ are allowed. The labels parallel ($\parallel$) and perpendicular ($\perp$) denote orientations relative to the magnetic field direction. The following derivations are for the plasma rest frame with cold plasma dispersion relation (Arons1986):
@@
The fields $\delta\mathbf{B}$ and $\delta\mathbf{E}$ of these waves in strong magnetic fields are polarized in the $\mathbf{k}-\mathbf{B}$ plane. In the chosen 1D configuration, Alfvén waves (AWs) do not propagate. They have no associated electric field, $\delta E_{y}=0$, the only non-vanishing component is $\delta B_{z}$. The static perturbations require a current $J_{y}=-(c/4\pi)\partial_{x}\delta B_{z}$ to fulfill Ampère-Maxwell’s law for $\partial_{t}E_{y}=0$ in the plasma rest frame.
@@ II.2 Frequency matching across the shock
Appendix A shows transformations of field and wave properties between relevant frames (u/s/d). In the shock front frame, wave frequencies match across the discontinuity due to energy conservation:
@@
Alfvénic perturbations with $\omega_{A}=0$ in their respective rest frames follow $\gamma_{\rm s}\beta_{\rm s}k_{\rm d}]^{\rm d}=[\gamma_{\rm s}\beta_{\rm s}k_{\rm u}]^{\rm u}$ (see Appendix A). For large upstream magnetizations and relativistic flow velocities, which is typical for strong shocks in magnetized plasmas, we find:
@@
For superluminal modes, frequency matching combined with Equation (3) implies a downstream wavenumber:
@@
Here, $\omega_{\rm pd}$ is the plasma frequency calculated with the downstream density, and we evaluated the dispersion relation in the shock front frame (Equation A5). When $\omega_{\rm u}]^{\rm s}<\omega_{\rm pd}$, no propagating modes exist. In the high-frequency limit $\omega_{\rm u}]^{\rm s}\gg\omega_{\rm pd}$, the downstream wavenumber can be approximated as (see Appendix A):
@@
The last approximation again requires high flow velocities in the shock-front frame. Equation (6) constrains the minimum wavenumber of seed perturbations to allow for propagating downstream modes:
@@
Here, $\omega_{\rm pu}$ is the plasma frequency calculated with the upstream density. We examine the scales for propagating and non-propagating waves across the shock normalized to the respective plasma skin depth and frequency. For the upstream wavenumber normalized by a plasma skin depth $d_{\rm u}]^{\rm d}=[d_{\rm u}]^{\rm u}/[\gamma_{\rm u}]^{\rm d}=c/(\omega_{\rm pu}[\gamma_{\rm u}]^{\rm d})$, we find:
@@ II.3 Amplitude matching across the shock
Non-propagating Alfvénic perturbations on a constant background $B_{0\rm u}$ in the upstream rest frame have electromagnetic (EM) fields
@@
MHD jump conditions (Appendix A.3) imply that electric fields are continuous in the frame of the shock. For relativistic shocks, the magnetic field experiences the compression $B_{y\rm d}]^{\rm d}/[B_{y\rm u}]^{\rm d}\approx 2$. Upstream waves at MHD scale will experience a similar field compression $\delta B_{z\rm d}]^{\rm d}/[\delta B_{z\rm u}]^{\rm d}\approx 2$. High-frequency waves with $\omega_{\rm u}]^{\rm d}\gtrsim\omega_{\rm pu}[\gamma_{\rm u}]^{\rm d}$ essentially propagate as EM waves with $\delta B_{z\rm d}]^{\rm d}/[\delta B_{z\rm u}]^{\rm d}\approx 1$ and continuous Poynting flux across the shock. Then, EM fields are continuous in all frames. In the downstream frame,
@@
For $c[k_{\rm d}]^{\rm d}\gg\omega_{\rm pd}$, the wave speed $\beta_{O}\approx-1$ (Equation 3), such that $\delta B_{A\rm d}]^{\rm d}\rightarrow 0$. At high-frequencies, only superluminal O-modes are generated downstream.
@@ III Simulations
We conduct 1D particle-in-cell (PIC) simulations of mode conversion at a magnetized relativistic perpendicular shock in pair plasma with the Tristan-MP.v2 code (tristanv2).
@@ III.1 Setup
Our simulation setup follows previous simulations of relativistic collisionless shocks (e.g., Spitkovsky2008; Sironi2009; Parsons2024). We initialize an upstream flow of magnetized pair plasma with Lorentz factor $\gamma_{\rm u}]^{d}$ along the $\mathbf{\hat{x}}$ direction and a thermal spread $T/m_{e}c^{2}=10^{-3}$ in units with $k_{B}=1$. The magnetic field is initially uniform with $\bm{B}_{0\rm u}]^{\rm d}=\left(B_{0\rm u}/B_{0}\right)\hat{\mathbf{y}}$. A conducting wall located at $x=0$ reflects particles and acts as a conducting boundary for fields. The counter-streaming flows form a shock that propagates along the $\mathbf{\hat{x}}$ direction. The simulation frame is the downstream rest frame. We simulate a domain with $N=96\times 10^{4}$ cells, and a skin-depth $d_{\rm u}]^{\rm d}=240$ cells. The total fixed-length domain extends for $N/[d_{\rm u}]^{\rm d}=4\times 10^{3}$ upstream plasma skin-depths. We choose an upstream magnetization $\sigma_{\rm u}]^{u}=[B_{0\rm u}^{2}/\gamma_{\rm u}]^{\rm d}$. The initial flow fills a region extending up to $x/[d_{\rm u}]^{\rm d}\lesssim 40$, fields and particles are continuously replenished by a moving injector located upstream of the shock. Once the shock has fully developed and propagates in the upstream plasma, the injector additionally generates wave fields in the simulation frame, as outlined below. Currents to support this perturbation are supplied by a particle velocity component along $\mathbf{\hat{y}}$ (see Section II.1). We track wave and shock dynamics for various properties of the upstream flow and upstream seed modes.
Figure 2: Shock and mode conversion dynamics of monochromatic seed waves. Left: Upstream scales $B_{0}=10$, $\gamma_{\rm u}]^{\rm d}=4$, $k_{\rm u}d_{\rm u}]^{\rm u}=0.63$. A superluminal O-mode (supO) propagates into the downstream, a small-amplitude non-propagating Alfvénic perturbation (AW) remains in the wake of the shock. The region of pure supO is shaded in blue, mixed supO/AW in red. The bottom panel shows EM fields associated with the propagating superluminal O-mode. Right: Upstream waves with $k_{\rm u}d_{\rm u}]^{\rm u}=0.03$. No modes propagate into the downstream, only AW perturbations remain (red-shaded region of top panels).

### v2

@@ Abstract
Fast radio bursts (FRBs) are extremely energetic radio transients, some are generated in magnetar magnetospheres and winds. Despite a growing number of observations, their emission mechanisms remain elusive. It has recently been proposed that Alfvénic perturbations can convert into superluminal O-modes at magnetized shocks and propagate downstream as a radio signal. We validate this superluminal wave activation mechanism using pair-plasma theory and particle-in-cell simulations. Theory predicts two different downstream modes: nonpropagating Alfvénic perturbations and propagating superluminal O-modes. Superluminal wave activation occurs if the frequency of upstream perturbations in the shock frame exceeds the downstream plasma frequency. 1D particle-in-cell simulations confirm wavenumber and frequency jumps across the shock for upstream perturbations with frequencies well above the plasma frequency. Our simulations model both monochromatic upstream waves and broadband spectra with the downstream plasma frequency acting like a high-pass filter for superluminal O-modes. We discuss implications for FRB generation in relativistic magnetized winds.
@@ I Introduction
Relativistic magnetized shocks generate many observable phenomena in extreme astrophysical environments, like particle acceleration to ultrahigh energies (e.g., Sironi and Spitkovsky, 2009; Sironi et al., 2015) and coherent wave generation via synchrotron maser emission (Iwamoto et al., 2017; Plotnikov and Sironi, 2019; Babul and Sironi, 2020; Sironi et al., 2021; Vanthieghem and Levinson, 2025a). Magnetars are neutron stars with strong magnetic field strengths $B\gtrsim 10^{14}\,{\rm G}$, slow rotation periods $P\sim 1\,{\rm s}$, and coronae of magnetically active plasmas (Rea and Esposito, 2010; Kaspi and Beloborodov, 2017). Their strongly magnetized and dynamic environments are a natural place for the formation of relativistic shocks at various scales (e.g., Margalit and Metzger, 2018; Metzger et al., 2019; Beloborodov, 2020; Thompson, 2022). Bursting, likely caused by crustal activity, can drive magnetospheric instabilities and seed magnetic fluctuations in expanding relativistic outflows.
@@
Thompson (2022) suggested that nonpropagating Alfvénic perturbations advected toward a magnetized shock convert to propagating superluminal kinetic plasma waves downstream. Such superluminal O-modes have frequencies $\omega/\bar{\omega}_{\rm p}\gtrsim 1$, where $\bar{\omega}_{\rm p}$ is the rest-frame plasma frequency (Arons and Barnard, 1986; Bransgrove et al., 2023). Superluminal O-modes could potentially escape compact object environments as radio emission, depending on the properties of the surrounding outflow.
@@
Transient radio activity, especially fast radio bursts (FRBs), are messengers of magnetically active environments both within (SGR 1935+2154; Andersen et al., 2020) and outside of our Galaxy (CHIME/FRB Collaboration et al., 2025). These powerful millisecond-duration flares are typically narrow band (a few hundred megahertz) and their peak frequency ranges between $120\,{\rm MHz}$ and several GHz (Pastor-Marazuela et al., 2021). They can carry information about their host environment, for example in polarization (e.g., Niu et al., 2024), and scintillation (Nimmo et al., 2025). FRBs are emerging as powerful probes of turbulent plasmas throughout the Universe (Connor et al., 2025; Ocker et al., 2025). However, their generation mechanisms remain elusive. This letter provides testable predictions of the Thompson (2022) model for radio wave generation at relativistic shocks in compact object magnetospheres. It is organized as follows. Section II outlines a simplified theoretical framework with relevant scales of the wave activation process. Section III shows 1D particle-in-cell (PIC) simulations (setup in Section III.1) of mode activation for monochromatic upstream fluctuations (III.2) and broadband seed wave spectra (III.3). We discuss results and their astrophysical implications in Section IV.
II One-dimensional Model of Alfvén Wave Activation
@@
Figure 1: Schematic visualization of mode conversion at relativistic magnetized shocks. A shock front propagates through a magnetized plasma. Nonpropagating Alfvénic perturbations upstream of the shock ($\omega_{A}=0$, waves in white) convert into a superposition of nonpropagating Alfvénic perturbations (white) and propagating superluminal O-modes (blue) in the downstream. Quantities measured in different frames are denoted as follows: downstream (d), upstream (u), shock (s).
@@
We consider a 1D magnetized planar shock that propagates along $\hat{\mathbf{x}}$ in a pair plasma with background magnetic field $\bm{B}_{0\rm u}]^{\rm d}=\left(B_{0\rm u}/B_{0}\right)\hat{\mathbf{y}}$ (see Figure 1). Here, $eB_{0}=m_{e}\omega_{Le0}c$ normalizes the magnetic field to the electron Larmor frequency $\omega_{Le0}$ in the downstream frame. Subscripts denote to which region the quantity belongs: the downstream (d), shock front (s), or upstream (u). Superscripts indicate in which frame a quantity is measured. This section reviews ordinary modes in cold 1D plasmas and describes a mode-conversion mechanism between subluminal (Alfvén) and superluminal modes at magnetized shocks (Figure 1). Nonpropagating Alfvénic perturbations frozen into the upstream flow are activated as superluminal modes propagating downstream.
II.1 One-dimensional Ordinary Modes in Cold Plasmas
@@
In 1D and with $\bm{B}_{0\rm u}$ perpendicular to $\hat{\mathbf{x}}$, only modes with wavenumber $k_{\parallel}=0$ are allowed. The labels parallel ($\parallel$) and perpendicular ($\perp$) denote orientations relative to the magnetic field direction. The following derivations are for the plasma rest-frame with cold plasma dispersion relation (Arons and Barnard, 1986):
@@
The fields $\delta\mathbf{B}$ and $\delta\mathbf{E}$ of these waves in strong magnetic fields are polarized in the $\mathbf{k}-\mathbf{B}$ plane. In the chosen 1D configuration, Alfvén waves (AWs) do not propagate. They appear as frozen-in magnetic shear fluctuations with no associated electric field, $\delta E_{y}=0$, and one nonvanishing component, $\delta B_{z}$. The static perturbations carry a current $J_{y}=-(c/4\pi)\partial_{x}\delta B_{z}$ to fulfill the Ampère-Maxwell law in the plasma rest-frame. We will refer to the transverse magnetic shear fluctuations with Alfvénic polarization as Alfvénic perturbations.
@@ II.2 Frequency Matching across the Shock
Appendix A shows transformations of field and wave properties between relevant frames (u, s, and d). In the shock front frame, wave frequencies match across the discontinuity due to energy conservation:
@@
Alfvénic perturbations with $\omega_{A}=0$ in their respective rest-frames follow $\gamma_{\rm s}\beta_{\rm s}k_{\rm d}]^{\rm d}=[\gamma_{\rm s}\beta_{\rm s}k_{\rm u}]^{\rm u}$ (see Appendix A). For large upstream magnetizations and relativistic flow velocities, which is typical for strong shocks in magnetized plasmas, we find
@@
For superluminal modes, frequency matching combined with Equation (3) implies a downstream wavenumber
@@
Here, $\omega_{\rm pd}$ is the plasma frequency calculated with the downstream density, and we evaluated the dispersion relation in the shock-front frame (Equation A5). When $\omega_{\rm u}]^{\rm s}<\omega_{\rm pd}$, no propagating modes exist. In the high-frequency limit $\omega_{\rm u}]^{\rm s}\gg\omega_{\rm pd}$, the downstream wavenumber can be approximated as (see Appendix A):
@@
The last approximation again requires high flow velocities in the shock front frame. Equation (6) constrains the minimum wavenumber of seed perturbations to allow for downstream-propagating modes:
@@
Here, $\omega_{\rm pu}$ is the plasma frequency calculated with the upstream density. We examine the scales for propagating and nonpropagating waves across the shock normalized to the respective plasma skin depth and frequency. For the upstream wavenumber normalized by a plasma skin depth $d_{\rm u}]^{\rm d}=[d_{\rm u}]^{\rm u}/[\gamma_{\rm u}]^{\rm d}=c/(\omega_{\rm pu}[\gamma_{\rm u}]^{\rm d})$, we find
@@ II.3 Amplitude Matching across the Shock
Nonpropagating Alfvénic perturbations on a constant background $B_{0\rm u}$ in the upstream rest-frame have electromagnetic (EM) fields
@@
MHD jump conditions (Appendix A.3) imply that electric fields are continuous in the frame of the shock. For relativistic shocks, the magnetic field experiences the compression $B_{y\rm d}]^{\rm d}/[B_{y\rm u}]^{\rm d}\approx 2$. Upstream waves at MHD scale will experience a similar field compression, $\delta B_{z\rm d}]^{\rm d}/[\delta B_{z\rm u}]^{\rm d}\approx 2$. High-frequency waves with $\omega_{\rm u}]^{\rm d}\gtrsim\omega_{\rm pu}[\gamma_{\rm u}]^{\rm d}$ essentially propagate as EM waves with $\delta B_{z\rm d}]^{\rm d}/[\delta B_{z\rm u}]^{\rm d}\approx 1$ and continuous Poynting flux across the shock. Then, EM fields are continuous in all frames. In the downstream frame
@@
For $c[k_{\rm d}]^{\rm d}\gg\omega_{\rm pd}$, the wave speed $\beta_{O}\approx-1$ (Equation 3), such that $\delta B_{A\rm d}]^{\rm d}\rightarrow 0$. At high frequencies, only superluminal O-modes are generated downstream.
@@ III Simulations
We conduct 1D PIC simulations of mode conversion at a magnetized relativistic perpendicular shock in pair plasma with the Tristan-MP.v2 code (Hakobyan and Spitkovsky, 2020).
@@ III.1 Setup
Our simulation setup follows previous simulations of relativistic collisionless shocks (e.g., Spitkovsky, 2008; Sironi and Spitkovsky, 2009; Parsons et al., 2024). We initialize an upstream flow of magnetized pair plasma with Lorentz factor $\gamma_{\rm u}]^{d}$ along the $\mathbf{\hat{x}}$ direction and a thermal spread $T/m_{e}c^{2}=10^{-3}$ in units with $k_{B}=1$. The magnetic field is initially uniform with $\bm{B}_{0\rm u}]^{\rm d}=\left(B_{0\rm u}/B_{0}\right)\hat{\mathbf{y}}$. A conducting wall located at $x=0$ reflects particles and acts as a conducting boundary for fields. The counterstreaming flows form a shock that propagates along the $\mathbf{\hat{x}}$ direction. The simulation frame is the downstream rest-frame. We simulate a domain with $N=96\times 10^{4}$ cells, and a skin depth $d_{\rm u}]^{\rm d}=240$ cells. The total fixed-length domain extends for $N/[d_{\rm u}]^{\rm d}=4\times 10^{3}$ upstream plasma skin depths. We choose an upstream magnetization $\sigma_{\rm u}]^{u}=[B_{0\rm u}^{2}/\gamma_{\rm u}]^{\rm d}$. The initial flow fills a region extending up to $x/[d_{\rm u}]^{\rm d}\lesssim 40$, and fields and particles are continuously replenished by a moving injector located upstream of the shock. Once the shock has fully developed and propagates in the upstream plasma, the injector additionally generates wave fields in the simulation frame, as outlined below. Currents to support this perturbation are supplied by a particle velocity component along $\mathbf{\hat{y}}$ (see Section II.1). We track wave and shock dynamics for various properties of the upstream flow and upstream seed modes.
Figure 2: Shock and mode-conversion dynamics of monochromatic seed waves. Left: upstream scales $B_{0}=10$, $\gamma_{\rm u}]^{\rm d}=4$, $k_{\rm u}d_{\rm u}]^{\rm u}=0.63$. A superluminal O-mode (supO) propagates downstream, a small-amplitude nonpropagating Alfvénic perturbation (AW) remains in the wake of the shock. The region of pure supO is shaded in blue, mixed supO/AW in red. The bottom panel shows EM fields associated with the propagating superluminal O-mode. Right: upstream waves with $k_{\rm u}d_{\rm u}]^{\rm u}=0.03$. No modes propagate downstream, only AW perturbations remain (red shaded region of top panels). We provide online animations of the wave activation dynamics in the supporting material (see Wotring, 2026a, b).
Figure 3: Wavenumber scaling (top) and frequency matching (bottom) across shocks with $\gamma_{\rm u}]^{\rm d}=4$ (left) and $\gamma_{\rm u}]^{\rm d}=8$ (right). Top panels show downstream wavenumbers as a function of upstream seed wavenumbers in the superposition region for AWs (red circles) and superluminal O-modes (blue circles). The scalings agree closely with Equations (10) and (11). Star markers denote cases for which no propagating downstream modes can be measured. Gray shading indicates the range between the plasma-frequency cutoff estimates derived in the cold plasma and warm plasma limits (using Equation 6). Bottom panels show the dispersion relation of propagating downstream modes. Downstream wavenumbers and corresponding wave frequencies trace the respective (hot) dispersion relation for the SupO-mode with temperatures $T_{\rm d}]^{\rm d}=0.96\,m_{e}c^{2}$ for $\gamma_{\rm u}]^{\rm d}=4$, and $T_{\rm d}]^{\rm d}=2.1\,m_{e}c^{2}$ for $\gamma_{\rm u}]^{\rm d}=8$ (see also Equation 18).